# Spread complexity in saddle-dominated scrambling

### v2

@@ 1 Introduction
Beyond its potential as a tool for probing quantum chaos, Krylov complexity has spurred extensive research due to its non-trivial connection with the spectral form factor and its implications for holography. [^4] The concept of Krylov complexity is gaining prominence, particularly in the context of holography, as it is well-defined in any quantum theory, overcoming the ambiguities in defining specific elementary gates or establishing tolerances. This interest has led to investigations across various domains, including research in diverse quantum systems Balasubramanian:2022tpr ; Barbon:2019wsy ; Avdoshkin:2019trj ; Dymarsky:2019elm ; Rabinovici:2020ryf ; Cao:2020zls ; Kim:2021okd ; Rabinovici:2021qqt ; Trigueros:2021rwj ; Fan:2022xaa ; Heveling:2022hth ; Bhattacharjee:2022vlt ; Caputa:2022eye ; Caputa:2022yju ; Muck:2022xfc ; Rabinovici:2022beu ; He:2022ryk ; Hornedal:2022pkc ; Alishahiha:2022anw ; Bhattacharjee:2022vlt ; Baek:2022pkt ; Bhattacharjee:2023dik ; Bhattacharyya:2023dhp ; Erdmenger:2023wjg ; Hashimoto:2023swv ; Camargo:2023eev ; Balasubramanian:2023kwd ; Bhattacharya:2023yec ; Bhattacharjee:2022qjw , gauge theories Magan:2020iac , holographic models Jian:2020qpp ; Rabinovici:2023yex , conformal/quantum field theories Dymarsky:2021bjq ; Caputa:2021ori ; Avdoshkin:2022xuw ; Camargo:2022rnt ; Adhikari:2022whf ; Vasli:2023syq , Lie groups Caputa:2021sib ; Patramanis:2021lkx ; Patramanis:2023cwz ; Chattopadhyay:2023fob , matrix models Iizuka:2023pov ; Iizuka:2023fba , models of quantum quenches Afrasiar:2022efk ; Pal:2023yik and open quantum systems Bhattacharya:2022gbz ; Liu:2022god ; Bhattacharjee:2022lzy ; Bhattacharya:2023zqt ; Bhattacharjee:2023uwx .
@@ 2.2 More on spread complexity: Ehrenfest theorem and early time growth
Additionally, it is also pertinent to note that the parameter $a/R$ serves as a relevant deformation parameter in the context of billiards BenettinStochastic ; McDonaldSpectrum ; CasatiSpectra systems. Notably, the circle billiard ($a/R=0$) exhibits characteristics typical of an integrable system, such as a vanishing Lyapunov exponent Hashimoto:2017oit , while the stadium billiard ($a/R\neq 0$) is regarded as non-integrable (i.e., chaotic systems), characterized by a finite Lyapunov exponent.
@@ 3.2.2 Spread complexity
In Fig. 9, we find that that the LMG model can indeed demonstrate the proposed resemblance between spread complexity and the SFF. In particular, our result suggests that the linear growth in spread complexity might be determined by the slope of SFF, which is not be exclusively tied to chaotic behavior. This observation aligns with the findings reported in Erdmenger:2023wjg .

### v3

@@ 1 Introduction
Beyond its potential as a tool for probing quantum chaos, Krylov complexity has spurred extensive research due to its non-trivial connection with the spectral form factor and its implications for holography. [^4] The concept of Krylov complexity is gaining prominence, particularly in the context of holography, as it is well-defined in any quantum theory, overcoming the ambiguities in defining specific elementary gates or establishing tolerances. This interest has led to investigations across various domains, including research in diverse quantum systems Balasubramanian:2022tpr ; Barbon:2019wsy ; Avdoshkin:2019trj ; Dymarsky:2019elm ; Rabinovici:2020ryf ; Cao:2020zls ; Kim:2021okd ; Rabinovici:2021qqt ; Trigueros:2021rwj ; Fan:2022xaa ; Heveling:2022hth ; Bhattacharjee:2022vlt ; Caputa:2022eye ; Caputa:2022yju ; Muck:2022xfc ; Rabinovici:2022beu ; He:2022ryk ; Hornedal:2022pkc ; Alishahiha:2022anw ; Baek:2022pkt ; Bhattacharjee:2023dik ; Bhattacharyya:2023dhp ; Erdmenger:2023wjg ; Hashimoto:2023swv ; Camargo:2023eev ; Balasubramanian:2023kwd ; Bhattacharya:2023yec ; Bhattacharjee:2022qjw ; Caputa:2024vrn , gauge theories Magan:2020iac , holographic models Jian:2020qpp ; Rabinovici:2023yex , conformal/quantum field theories Dymarsky:2021bjq ; Caputa:2021ori ; Avdoshkin:2022xuw ; Camargo:2022rnt ; Adhikari:2022whf ; Vasli:2023syq , Lie groups Caputa:2021sib ; Patramanis:2021lkx ; Patramanis:2023cwz ; Chattopadhyay:2023fob , matrix models Iizuka:2023pov ; Iizuka:2023fba , models of quantum quenches Afrasiar:2022efk ; Pal:2023yik and open quantum systems Bhattacharya:2022gbz ; Liu:2022god ; Bhattacharjee:2022lzy ; Bhattacharya:2023zqt ; Bhattacharjee:2023uwx .
@@
Furthermore, it is noteworthy that a recent study Caputa:2024vrn provides a thorough examination of the comparison and contrast between Krylov operator complexity and spread complexity. Their findings reveal that in general both complexities can exhibit disparate behaviors, particularly in the intermediate time regime where the peak of spread complexity becomes manifest. Our work thus can stand as a distinct and original contribution to the field, separate from the analyses of Krylov operator complexity conducted in Bhattacharjee:2022vlt ; Baek:2022pkt .
@@ 2.2 More on spread complexity: Ehrenfest theorem and early time growth
Additionally, it is also pertinent to note that the parameter $a/R$ serves as a relevant deformation parameter in the context of billiards Benettin:1978aa ; McDonaldSpectrum ; Casati1980 systems. Notably, the circle billiard ($a/R=0$) exhibits characteristics typical of an integrable system, such as a vanishing Lyapunov exponent Hashimoto:2017oit , while the stadium billiard ($a/R\neq 0$) is regarded as non-integrable (i.e., chaotic systems), characterized by a finite Lyapunov exponent.
@@ 3.2.2 Spread complexity
In Fig. 9, we find that the LMG model can indeed demonstrate the proposed resemblance between spread complexity and the SFF. In particular, our result suggests that the linear growth in spread complexity might be determined by the slope of SFF, which is not be exclusively tied to chaotic behavior. This observation aligns with the findings reported in Erdmenger:2023wjg .
@@
It is worth noting that the diminishment of the characteristic feature of quantum choas at lower temperatures (i.e., finite $\beta$) may not be a surprise. Analogous phenomena have been observed in related studies, such as the transition of the exponential growth of out-of-time-ordered correlators (OTOCs) to oscillatory behavior with decreasing temperature: e.g., see Fig. 13 in Akutagawa:2020qbj .